calculate the bond length of C-X bond if C-C bond length is 1.54 A* and X-X bond length is 1.2 A* and electronegativities of C & X are 2.0 & 3.0 respectively.

Grade:11

4 Answers

ankit agrawal
12 years ago

rc= dc-c/2 = 1.54/2 = 0.77

rx = dx-x/2 = 1.2/2 = 0.6

dc-x = rc + rx - 0.09(ex-ec)

      = 0.77 + 0.6 - 0.09(3-2)

      = 1.46 A*    

Pralabh
7 years ago
Can you please explain the formula you used. I mean from where did you get dc-x=rc+rx-0.09(ex-ec) etc. And how is it derived?
Bunny
6 years ago
The mathematical formula is absolutely right.but my dear brother the calculation is wrong.the ans will be 1.28 A°
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

rc= dc-c/2 = 1.54/2 = 0.77
rx= dx-x/2 = 1.2/2 = 0.6
dc-x= rc + rx- 0.09(ex-ec)
= 0.77 + 0.6 - 0.09(3-2)
= 1.28 A*

Thanks and Regards